# Modified Dai-Liao Spectral Conjugate Gradient Method with Application to Signal Processing

### v2

@@ Abstract
In this article, we present a modified variant of the Dai-Liao spectral conjugate gradient method, developed through an analysis of eigenvalues and inspired by a modified secant condition. We show that the proposed method is globally convergent for general nonlinear functions under standard assumptions. By incorporating the new secant condition and a quasi-Newton direction, we introduce updated spectral parameters. These changes ensure that the resulting search direction satisfies the sufficient descent property without relying on any line search. Numerical experiments show that the proposed algorithm performs better than several existing methods in terms of convergence speed and computational efficiency. Its effectiveness is further demonstrated through an application in signal processing.
@@
Spectral Conjugate Gradient Method , Dai-Liao Method , Optimization , Signal Processing
MSC:
[2020] 65K05 , 90C30 , 49M37
@@ 1 Introduction
[…]
find $\text{find }x^{*}\in\mathbb{R}^{m}\text{ such that }f(x^{*})=\min_{x\in\mathbb{R}^{m}}f(x).$ (1)
We denote the set of minimizers of $f$ as $\Omega_{\text{min}}(f,\mathbb{R}^{m})$.
The general iterative rule for solving (1) starts from an initial point $x_{0}\in\mathbb{R}^{m}$ and generates a sequence $\{x_{n}\}$ in $\mathbb{R}^{m}$ as follows:
@@
where the stepsize $\alpha_{n}$ is a positive real parameter determined via exact or inexact line search, $x_{n}$ is the previous iterative point, $x_{n+1}$ is the current point and $d_{n}$ is a suitable search direction. The class of algorithms that follow the form (2) is known as line search algorithms. For such algorithms, only the search direction $d_{n}\in\mathbb{R}^{m}$ and the stepsize $\alpha_{n}\in(0,\infty)$ are required. An appropriate descent direction $d_{n}$ must be chosen such that it satisfies the descent condition:
@@
A common choice for a descent direction is $d_{n}=-\nabla f(x_{n})$, which reduces the general line search iteration (2) to the gradient descent (GD) method:
@@
Conjugate gradient (CG) methods are among the most effective algorithms for solving large-scale unconstrained optimization problems due to their low storage requirements. These methods require only first-order derivatives, making them highly suitable for large-scale applications. The first form of the CG method was introduced by Hestenes and Stiefel [1] in 1952 for solving linear systems of equations. Later, in 1964, Fletcher and Reeves [2] modified the original idea and developed the first nonlinear CG method for unconstrained optimization problems. Due to the strong theoretical foundation and excellent numerical performance of these algorithms, many variants of the CG method have been proposed (see, e.g., [3, 4, 5, 6]).
The successive approximations generated by the CG algorithms have the form (2), where $\alpha_{n}\in(0,\infty)$ is the step length determined through a line search along the CG direction $d_{n}$. The CG directions $d_{n}$ is defined as:
@@
where $\beta_{n}$ is the CG parameter. There are many formulas for $\beta_{n}$, each resulting in different computational behavior [7].
Based on an extended conjugacy condition, one of the fundamental CG methods was proposed by Dai and Liao [4] with the CG parameter defined as:
@@
where $t\in(0,\infty)$, $s_{n}=x_{n+1}-x_{n}$, and $y_{n}=\nabla f(x_{n+1})-\nabla f(x_{n})$. Note that when $t=0$, $\beta_{n}^{DL}$ simplifies to the CG parameter proposed by Hestenes and Stiefel [1]:
@@
The CG parameter proposed by Hager and Zhang [8] is defined as:
$\beta_{n}^{HZ}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$
which can be viewed as an adaptive version of (5) for the specific choice $t=2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}$. Similarly, the CG parameter proposed by Dai and Kou [9] is defined as:
@@
where $\tau_{n}$ is a parameter corresponding to the scaling factor in the scaled memoryless Broyden-Fletcher-Goldfarb-Shanno (BFGS) method, which can be regarded as another adaptive form of (5).
Inspired by Shanno’s matrix perspective on CG algorithms [10], Babaie-Kafaki and Ghanbari [11] observed that the Dai-Liao search directions can be expressed as $d_{n+1}=-P_{n+1}\nabla f(x_{n+1})$ for all $n\in\mathbb{N}_{0}$, where $P_{n+1}=I-\frac{s_{n}y_{n}^{T}}{\langle s_{n},y_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}$ is a nonsingular matrix for $t\in(0,\infty)$ and $\langle s_{n},y_{n}\rangle\neq 0$; see [12]. Conducting an eigenvalue analysis on a symmetrized version of $P_{n+1}$, defined by $A_{n+1}=\frac{P_{n+1}+P_{n+1}^{T}}{2}$, we can derive a two-parameter formula for $t$ in the Dai-Liao CG parameter (5), given by
@@
which ensures the descent condition (3) for $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. However, in cases where $\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}$ and $\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$ are nearly equal, their subtraction may not be numerically stable [13]. Moreover, the parameter $t$ in CG parameter $\beta_{n}^{DL}$ must be nonnegative. To address these issues, they restricted $q$ in (6) to be nonpositive and proposed the following CG parameter:
@@
where $t_{n}^{p,q}$ is defined by (6) with $p\in\left(\frac{1}{4},\infty\right)$ and $q\in(-\infty,0]$. Note that for the parameter choice $(p,q)=(2,0)$ in (6), $\beta_{n}^{DDL}$ reduces to $\beta_{n}^{HZ}$. Similarly, if $(p,q)=(1,0)$, then $\beta_{n}^{DDL}$ is equivalent to $\beta_{n}^{DK}(\tau_{n})$ with the optimal choice of $\tau_{n}=\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$.
Another significant category of iterative methods used for solving unconstrained optimization problems is the spectral gradient method. Initially introduced by Barzilai and Borwein [14], these methods were later advanced by Raydan [15], who developed a spectral gradient approach specifically for large-scale unconstrained optimization challenges. The spectral conjugate direction is defined by
@@
where $\{\theta_{n}\}\subset(0,\infty)$ is spectral parameter and $\{\beta_{n}\}\subset(0,\infty)$ is conjugate parameter.
Building on the encouraging numerical results associated with spectral gradient methods [14, 15], Birgin and Martínez [16] proposed the integration of spectral gradient techniques with conjugate gradient methods, culminating in the development of the first spectral conjugate gradient (CG) method. The spectral CG methods represent a powerful enhancement of traditional CG methods by incorporating the principles of spectral gradients within the conjugate gradient framework. Furthering this line of research, Yu et al. [17] modified the spectral Perry’s CG method [18], resulting in an efficient spectral CG variant. Capitalizing on the robust convergence properties of quasi-Newton methods, Andrei [19] introduced a scaled BFGS preconditioned CG method in 2007, leveraging quasi-Newton techniques to enhance the effectiveness of conjugate gradient methods. This foundational idea has been elaborated upon by Andrei [20, 21, 22], as well as Babaie-Kafaki and Mahdavi-Amiri [23].
Using the modified secant condition proposed by Li and Fukushima [24], Zhou and Zhang [25] established the following modified secant condition:
$\nabla^{2}f(x_{n+1})s_{n}=z_{n},$
@@
with $r,\nu\in(0,\infty)$. The modified secant condition (7) plays a crucial role in establishing the global convergence of the MBFGS method for non-convex functions; see [25].
It is worth noting that the following relation holds independently of the line search:
@@
Inspired from the appropriate theoretical properties given by (7), Faramarzi and Amini [26] modified the spectral CG method proposed by Jian et al. [27] and considered the following conjugate parameter
@@
Based on this conjugate parameter, they construct the following spectral parameter:
@@
where $\eta$ and $\tau$ are positive constants. To ensure the global convergence of this method, Faramarzi and Amini [26] used the line search satisfying the strong Wolfe conditions:
@@
where $10<\delta<\sigma<$ 1 and the following assumptions:
@@
    The level set
    $C_{0}=\{x\in\mathbb{R}^{n}:f(x)\leq f(x_{0})\}$
    is bounded; namely, there exists a constant $M>0$ such that
    $\|x\|\leq M\text{ for all }x\in C_{0}.$
@@
    In some neighbourhood $\Omega$ of $C_{0}$, $f$ is continuously differentiable and its gradient is Lipschitz continuous, namely there exists a constant $L>0$ such that
    $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|\text{ for all }x,y\in\Omega.$ (10)
The primary motivation of this study is to introduce a modified spectral conjugate gradient method and to establish its global convergence for general functions under standard assumptions. This new method consistently satisfies the sufficient descent property, regardless of the line search used. To demonstrate the effectiveness of the proposed approach, we present numerical results that validate its performance and show its superior convergence compared to the modified spectral conjugate gradient method of Faramarzi and Amini [26] and the scaled conjugate gradient method of Mard and Fakhari [28]. Furthermore, we explore its application in compressed sensing, emphasizing its practical relevance and significant advantages.
@@ 2 Preliminaries
In this section, we present some fundamental results from Euclidean space that will be used to establish the convergence of the proposed method.
@@
([26]) Let $a,b,\sigma\in\mathbb{R}$. Then
$(c+\sigma d)^{2}\leq(1+\sigma^{2})(c^{2}+d^{2})\text{ for all }c,d,\sigma\geq 0,$
@@
([26]) Let $\mathbb{R}^{m}$ be the Euclidean space. Then
@@
Let $\mathbb{R}^{m}$ be the Euclidean space and let $\{u_{n}\}$ and $\{v_{n}\}$ be sequences in $\mathbb{R}^{m}$. Let $b_{0},\sigma\in(0,\infty)$ and $u_{0}\in\mathbb{R}^{m}$ and let $\{a_{n}\}$ and $\{b_{n}\}$ be sequences in $(0,\infty)$ satisfying the following conditions:
@@
    $\langle v_{n+1},u_{n+1}\rangle<0$ for all $n\in\mathbb{N}$,
@@
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}+a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\geq\frac{a}{1+\sigma^{2}}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$
@@
From (11) and condition (C2), we obtain
@@
$\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\geq b_{n}^{2}-a^{2}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$ (12)
Again, from (11), we have
$-\langle v_{n+1},u_{n+1}\rangle+b_{n}\langle v_{n+1},u_{n}\rangle=a_{n+1}\|v_{n+1}\|^{2}.$ (13)
From condition (C3), we have
@@
From (13) and (14), we get
$-\langle v_{n+1},u_{n+1}\rangle+\sigma b_{n}|\langle u_{n},v_{n}\rangle|\geq a_{n+1}\|v_{n+1}\|^{2}.$
@@
Since $a_{n+1}>a$, this yields that
$\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}$ $\geq\frac{a}{1+\sigma^{2}}\|v_{n+1}\|^{4},$
@@
$\geq\frac{1}{\|u_{n+1}\|^{2}}\left[\frac{a}{1+\sigma^{2}}\|v_{n+1}\|^{4}+\left(\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}-b_{n}^{2}\right)\langle v_{n},u_{n}\rangle^{2}\right].$ (15)
From (12) and (15), we conclude that
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n+1},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}\geq\frac{\|v_{n+1}\|^{4}}{\|u_{n+1}\|^{2}}\left[\frac{a}{1+\sigma^{2}}-a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\right],$
which implies that
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}+a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\geq\frac{a}{1+\sigma^{2}}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$
@@
([29]) Let $f:\mathbb{R}^{n}\to\mathbb{R}$ such that assumptions (A1) and (A2) hold. Consider any iterative method of the form (2), where $d_{n}$ is a descent direction and $\alpha_{n}$ is obtained by the strong Wolfe conditions (8) and (9). Then
@@
From Lemma 4, we have

### v3

# A Modified Dai-Liao Spectral Conjugate Gradient Method with an Application to Signal Processing

@@ Abstract
We propose and study a variant of the Dai-Liao spectral conjugate gradient method, developed through an analysis of eigenvalues and inspired by a modified secant condition. We show that our proposed method is globally convergent for general nonlinear functions under standard assumptions. By incorporating the new secant condition and a quasi-Newton direction, we introduce updated spectral parameters. These changes ensure that the resulting search direction satisfies the sufficient descent property without relying on any line search. Numerical experiments show that the proposed algorithm performs better than several existing methods in terms of convergence speed and computational efficiency. Its effectiveness is further demonstrated through an application to signal processing.
@@
Dai-Liao Method , Optimization , Signal Processing , Spectral Conjugate Gradient Method
2020 MSC:
65K05 , 90C30 , 49M37
@@
organization=Department of Mathematics, The Technion – Israel Institute of Technology,city=Haifa, country=Israel
@@ 1 Introduction
Let $\mathbb{R}^{m}$ denote the Euclidean space equipped with the usual inner product $\langle\cdot,\cdot\rangle$ and its associated norm $\|\cdot\|$. Let $f:\mathbb{R}^{m}\to\mathbb{R}$. Our goal is to solve the following minimization problem:
$\text{find }x^{*}\in\mathbb{R}^{m}\text{ such that }f(x^{*})=\min_{x\in\mathbb{R}^{m}}f(x).$ (1)
We denote the set of minimizers of $f$ by $\Omega_{\text{min}}(f,\mathbb{R}^{m})$.
A common approach to find a solution to (1) is the following iterative method: Let $x_{0}\in\mathbb{R}^{m}$ be an initial point. Given the current iterate $x_{n}$, the next iterate $x_{n+1}$ is computed by
@@
where $\alpha_{n}>0$ is the stepsize determined by exact or inexact line search and $d_{n}$ is a suitable search direction. Methods that follow the update rule (2) are referred to as line search methods. These methods require only the search direction $d_{n}\in\mathbb{R}^{m}$ and the stepsize $\alpha_{n}\in(0,\infty)$. An appropriate descent direction $d_{n}$ must be chosen so that it satisfies the descent condition:
@@
A common choice for a descent direction is $d_{n}=-\nabla f(x_{n})$, which simplifies the general line search method (2) to the familiar gradient descent (GD) method:
@@
Conjugate gradient (CG) methods are among the most effective algorithms for solving large-scale unconstrained optimization problems due to their low storage requirements. These methods require only first-order derivatives, makes them highly suitable for large-scale applications. The conjugate gradient (CG) method was originally proposed by Hestenes and Stiefel [Hestenes1952] for the solution of linear systems. Later, Fletcher and Reeves [Fletcher] extended this concept and formulated the first nonlinear CG method for unconstrained optimization. Due to their strong theoretical foundation and excellent numerical performance, many other versions of the CG method have been proposed (see [Polak1969, Dai2001, Fletcher1987, Dai1999] and the references therein).
In general, CG methods use the iterative scheme (2), where the search direction $d_{n}$ is defined by
@@
where $\beta_{n}$ is the CG parameter. There are many formulas for $\beta_{n}$, each resulting in different computational behavior [HagerZhang2006].
One of the most well-known CG methods was proposed by Dai and Liao [Dai2001]. Their approach is based on an extended conjugacy condition, where the CG parameter is defined by
@@
where $t\in(0,\infty)$, $s_{n}=x_{n+1}-x_{n}$, and $y_{n}=\nabla f(x_{n+1})-\nabla f(x_{n})$. Note that when $t=0$, $\beta_{n}^{DL}$ simplifies to the CG parameter proposed by Hestenes and Stiefel [Hestenes1952]:
@@
The CG parameter proposed by Hager and Zhang [HagerZhang2005] is defined by
$\beta_{n}^{HZ}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
The parameter $\beta_{n}^{HZ}$ can be viewed as an adaptive version of (5) for the specific choice $t=2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}$. Similarly, the CG parameter proposed by Dai and Kou [DaiKou2013] is defined by
@@
where $\tau_{n}$ is a parameter that plays the role of a scaling factor in the scaled memoryless Broyden–Fletcher–Goldfarb–Shanno (BFGS) method, which can be viewed as another adaptive variant of
Motivated by Shanno’s matrix interpretation of CG methods [Shanno1978], Babaie-Kafaki and Ghanbari [KafakiGh2014] showed that the Dai–Liao search directions can be written in the form $d_{n+1}=-P_{n+1}\nabla f(x_{n+1})$ for all $n\in\mathbb{N}_{0}$, where $P_{n+1}=I-\frac{s_{n}y_{n}^{T}}{\langle s_{n},y_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}$ is a nonsingular matrix for $t\in(0,\infty)$ and $\langle s_{n},y_{n}\rangle\neq 0$; see also [KafakiEJOR2014]. By performing an eigenvalue analysis on the symmetrized version of $P_{n+1}$, defined by $A_{n+1}=\frac{P_{n+1}+P_{n+1}^{T}}{2}$, we can derive a two-parameter formula for $t$ in the Dai-Liao CG parameter (5), given by
@@
which ensures the descent condition (3) for $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. However, in cases where $\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}$ and $\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$ are nearly equal, their subtraction may not be numerically stable [Kafaki2016]. Moreover, the parameter $t$ in the CG parameter $\beta_{n}^{DL}$ must be nonnegative. To address these issues, they restricted $q$ in (6) to be nonpositive and proposed the following CG parameter:
@@
where $t_{n}^{p,q}$ is defined by (6) with $p\in\left(\frac{1}{4},\infty\right)$ and $q\in(-\infty,0]$. Note that for the parameter choice $(p,q)=(2,0)$ in (6), $\beta_{n}^{DDL}$ reduces to $\beta_{n}^{HZ}$. Similarly, if $(p,q)=(1,0)$, then $\beta_{n}^{DDL}$ is equivalent to $\beta_{n}^{DK}(\tau_{n})$ with the optimal choice of $\tau_{n}=\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$. For additional variants of the Dial–Liao conjugate gradient method, we refer to [Kafaki2025, Kafaki2025Opt].
Another significant category of iterative methods used for solving unconstrained optimization problems is the spectral gradient method. Initially introduced by Barzilai and Borwein [Barzilai1988], these methods were later advanced by Raydan [Raydon1997], who developed a spectral gradient approach specifically for large-scale unconstrained optimization challenges. The spectral conjugate direction is defined by
@@
where $\{\theta_{n}\}\subset(0,\infty)$ is the spectral parameter and $\{\beta_{n}\}\subset(0,\infty)$ is the conjugate parameter.
Building on the encouraging numerical results associated with spectral gradient methods [Barzilai1988, Raydon1997], Birgin and Martínez [Birgin2001] proposed the integration of spectral gradient techniques with conjugate gradient methods, culminating in the development of the first spectral conjugate gradient (CG) method. The spectral CG methods represent a powerful enhancement of traditional CG methods by incorporating the principles of spectral gradients within the conjugate gradient framework. Continuing in this direction, Yu and co-authors [Yu2008OMS] adapted Perry’s spectral CG formulation [Perry1978], obtaining a more effective variant of the method. Inspired by the good convergence of quasi-Newton techniques, Andrei [Andrei2007] proposed a CG method using a scaled BFGS preconditioner, and this idea was further developed in later works [Andrei2008, Andrei2010EJOR, Babaie2013MMA].
Li and Fukushima [Li2001] proposed a modified form of the secant condition, and Zhou and Zhang [Zhou2006] later built on their idea to obtain the following modified secant relation:
$\nabla^{2}f(x_{n+1})s_{n}=z_{n},$ (7)
@@
with $r,\nu\in(0,\infty)$. The modified secant condition (8) is essential to prove the global convergence of the MBFGS algorithm when dealing with non-convex objectives [Zhou2006]. Note that the following expression holds irrespective of the line search used.
@@
Since (8) leads to favorable theoretical properties, Faramarzi and Amini [Faramarzi2019] modified the spectral CG method of Jian et al. [Jian2017OMS] and proposed the following choice for the conjugate parameter:
@@
Based on this conjugate parameter, they constructed the following spectral parameter:
@@
where $\eta$ and $\tau$ are positive constants. To ensure global convergence of this method, Faramarzi and Amini [Faramarzi2019] used the line search that satisfied the strong Wolfe conditions
@@
where $0<\delta<\sigma<1$, and the following assumptions:
@@
    The level set $C_{0}=\{x\in\mathbb{R}^{n}:f(x)\leq f(x_{0})$ is bounded.
@@
    $f$ is continuously differentiable in some neighbourhood $\Omega$ of $C_{0}$ with $L$-Lipschitz continuous gradient.
The primary motivation of the present study is to propose a modified spectral conjugate gradient method, guided by an eigenvalue-based analysis and a modified secant condition, and to prove global convergence for general objective functions under typical conditions.This new method always meets the sufficient-descent requirement, independent of the line search strategy applied. To demonstrate the effectiveness of our proposed approach, we present numerical results that validate its performance and show its superior convergence compared to the modified spectral conjugate gradient method of Faramarzi and Amini [Faramarzi2019], and the scaled conjugate gradient method of Mard and Fakhari [Mrad2024]. Furthermore, we explore its application to compressed sensing, emphasizing its practical relevance and significant advantages.
@@ 2 Preliminaries
In this section we present some fundamental results regarding Euclidean space that will be used to establish the convergence of our proposed method.
@@
([Faramarzi2019]) Let $a,b,\sigma\in\mathbb{R}$. Then
$(c+\sigma d)^{2}\leq(1+\sigma^{2})(c^{2}+d^{2})\text{ for all }c,d,\sigma\geq 0.$
@@
([Faramarzi2019]) Let $\mathbb{R}^{m}$ be a Euclidean space. Then
@@
Let $\mathbb{R}^{m}$ be a Euclidean space, and let $\{u_{n}\}$ and $\{v_{n}\}$ be sequences in $\mathbb{R}^{m}$. Let $\sigma\in(0,\infty)$ and let $\{a_{n}\}$ and $\{b_{n}\}$ be two sequences in $(0,\infty)$ satisfying the following conditions:
@@
    $\langle v_{n},u_{n}\rangle<0$ for all $n\in\mathbb{N}$,
@@
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}\geq\frac{\|v_{n+1}\|^{4}}{\|u_{n+1}\|^{2}}\left[\frac{a^{2}}{1+\sigma^{2}}-a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\right].$
@@
Using (11) and condition (C2), we obtain
@@
$\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\geq b_{n}^{2}-a_{n+1}^{2}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$ (12)
Again, using (11), we have
$\langle v_{n+1},u_{n+1}\rangle+b_{n}\langle v_{n+1},u_{n}\rangle=a_{n+1}\|v_{n+1}\|^{2}.$ (13)
It follows from condition (C3) that
@@
Using (13) and (14), we get
$\langle v_{n+1},u_{n+1}\rangle+\sigma b_{n}|\langle u_{n},v_{n}\rangle|\geq a_{n+1}\|v_{n+1}\|^{2}.$
@@
Since $a_{n+1}>a$, it follows that
$\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}$ $\geq\frac{a^{2}}{1+\sigma^{2}}\|v_{n+1}\|^{4},$
@@
$\geq\frac{1}{\|u_{n+1}\|^{2}}\left[\frac{a^{2}}{1+\sigma^{2}}\|v_{n+1}\|^{4}+\left(\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}-b_{n}^{2}\right)\langle v_{n},u_{n}\rangle^{2}\right].$ (15)
Using (12) and (15), we conclude that
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}\geq\frac{\|v_{n+1}\|^{4}}{\|u_{n+1}\|^{2}}\left[\frac{a^{2}}{1+\sigma^{2}}-a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\right].$
@@
([Zoutendijk]) Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be such that assumptions (A1) and (A2) hold. For any iterative method of the form (2) with $d_{n}$ being a descent direction and $\alpha_{n}$ determined by the strong Wolfe conditions (9) and (10), we have
@@
It follows from Lemma 4 that